The equilibrium constant for the following reaction

Ag_{2}O(s)\rightarrow 2Ag(s)+\frac{1}{2}O_{2}(g)

can be expressed by the empirical formula

\ln K_{P}=0.9692+\frac{5612.7K}{T}-\frac{2.0953\times 10^6K^2}{T^2}

in the interval 445 K < T < 460 K. Use this expression to calculate ∆rH◦ and ∆rS ◦ at 450 K and derive an equation for ∆fH◦ of Ag2O (s) in this temperature range. Now use the following heat capacity data:

C^{^{\circ}}_{P}[O_{2}(g)]/R=3.27+(5.03\times 10^{-4}K^{-1})T

C^{^{\circ}}_{P}[Ag(s)]/R=2.82+(7.55\times 10^{-4}K^{-1})T

C^{^{\circ}}_{P}[Ag_{2}O(s)]/R=6.98+(4.48\times 10^{-3}K^{-1})T

to calculate ∆rH◦ , ∆rS◦ , and ∆rG◦ at 298 K. Would Ag2O be automatically decomposed when exposed to the atmosphere (the partial pressure of O2 in the atmosphere is about 0.2 atm)?
Answer #1

Pas. .. Aggo (s) + 2Ag(s) + £ Q2 Ca) Given thals In Kp = 0.9692 + 5612.7K _ 2.0953x109K T temperature range, 445K<<< 460k #AHog Afisa - 30.76 kJ/mal 4 50K + RX 0.09692 + 5612.7 T - 2.0953710 (450) Os, Ars = 0.068 kJ[x]mal + 8.314 ( 212234 ) J.xmalNow, Ti = 450k, Am Husok = 30.46 K3/med. T2 = 298k ACP = 2*(CP) Agco + 5* ().gro) - (CP) Agora) on, AC = T2+{282 + (7.85x104)K T T D # Ange at 298k APG y = – RT In Kp e branz-RT (0,9692 + 5612. * - 3:09.534164) putting the value, T = 298k , we get, A